OpenStudy (anonymous):

HELP! SO CLOSE TO FINISH!!! What is the sum of the geometric sequence -3, 18, -108, ... if there are 8 terms?

OpenStudy (anonymous):

please explain!!!!!!!

OpenStudy (anonymous):

help?????

OpenStudy (ankit042):

can you find the common ratio of this series?

OpenStudy (anonymous):

the common ratio r is -6 and a1=-3 so the sum will be if there are 8 terms s8=-3((1-(-6)^8)/1-(-6)=143967

OpenStudy (anonymous):

719,835 119,973 -119,973 -719,835 these are the answer choices..

OpenStudy (anonymous):

its 719,835 i made a calc eror

OpenStudy (anonymous):

can you show me how?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sum=(a(1-r^n)/(1-r) a=-3 r=-6 n=8 sum=(-3(1-(-6)^8)/(1-(-6))=(-3*(1-1679616))/7=(-3*(-1679615))/7=5038845/7=719835.

OpenStudy (anonymous):

Sn=a1(1-r^n)/(1-r) you need to find the a1 the first term in your sequence which is -3 , and the common ratio, then you subst their values in the equation above , so the common ratio is -6 because if we multiply -3 by 6 we get 18 and if we multiply 18 by -6 we get -108 , you want to find the 8 terms so n=8 .